OpenStudy (anonymous):

I GIVE MEDALS! What is the equation of the line that passes through (1, 2) and is parallel to the line whose equation is 4x + y - 1 = 0? 4x - y - 6 = 0 4x + y - 6 = 0 4x + y + 6 = 0

OpenStudy (shamil98):

which one of those equations when evaluated at (1,2) make the equation true?..

OpenStudy (anonymous):

the last one?

OpenStudy (shamil98):

4(1) + 2 + 6 = 0 really? ..

OpenStudy (anonymous):

im really bad at this okay

OpenStudy (shamil98):

try again

OpenStudy (anonymous):

A

OpenStudy (shamil98):

Stop guessing -_- and no

OpenStudy (anonymous):

so its b

OpenStudy (shamil98):

yes.. (1,2) 4x + y - 6 = 0 4(1) + 2 - 6 = 0 6-6 =0 it checks out so yes